Need help please with solving equations involving rational exponents.
Solve equation. Check answer.
2x^(1/3) - 5 = 1

Answer by sudhanshu_kmr(1152)


2x^(1/3) - 5 = 1
=> 2x^(1/3) = 6
=> x^(1/3) = 6/2
=> x^(1/3) = 3
=> x = 3^3 = 27

x= 27